# Connecting the Dots: Is Mode-Connectedness the Key to Feasible Sample-Based Inference in Bayesian Neural Networks?

### v1

@@ C.3 Practical SBI
C.4 Dying Sampler and Bayesian Deep Ensembles
BDEs reported in Table 3 use 12 chains each with 8,000 samples (4,000 for the larger data sets) using the NUTS sampler with a short warmup phase of 100 steps. The table reports the first 10, 100 and 1,000 samples from these chains. Each chain’s initial proposal is a parameter set of the converged DE member trained on the same architecture, which is a network with two hidden layers of 16 neurons each. The priors on the weights are again unit Gaussians, leading to a valid initialization of the chains. The activation is ReLU.
@@ Appendix D Experimental Settings and Further Details
Our software is implemented in Python and mainly relies on the jax (Bradbury et al., 2018), numpyro (Phan et al., 2019) and pytorch (Paszke et al., 2019) libraries. We further use Docker for a reproducible experimental setup. Our code is available at https://anonymous.4open.science/r/bnn_connecting_the_dots-E2DF/.
@@
The experiments were run on 4 CPU instances with 32 cores each and 64GB RAM. Sampling 12 chains for most experiments allowed to parallelize the sampling such that at all times 2 experiments with 12 chains each can be run. The sampling with NUTS of 12 chains, 10,000 warmup steps and 8,000 samples (4,000 for the larger data sets) required on average three hours for the smaller data sets, and 30 hours for the larger ones on a two-hidden-layer network with 16 neurons in each hidden layer and tanh activation. This results in about half a second and 8 seconds for 12 parallel samples for the small and large data sets, respectively. Decreasing the number of warmup samples as proposed in BDEs reduces the number of required samples per chain by more than half, resulting in much faster model fits.

### v2

@@ C.3 Practical SBI
C.4 Dying Sampler and Deep Ensemble Initialization
DEI-MCMC reported in Table 3 uses 12 chains each with 8,000 samples (4,000 for the larger data sets) using the NUTS sampler with a short warmup phase of 100 steps. The table reports the first 10, 100 and 1,000 samples from these chains. Each chain’s initial proposal is a parameter set of the converged DE member trained on the same architecture, which is a network with two hidden layers of 16 neurons each. The priors on the weights are again unit Gaussians, leading to a valid initialization of the chains. The activation is ReLU.
Moreover, the comparison presented in Table 8 of the LPPD between DEI-MCMC and Deep Ensembles indicates a significant enhancement in uncertainty quantification also for deeper architectures.
Table 8: Average LPPD values (± their standard deviations across replications) of a DE, and DEI-MCMC (i.e., a BNN with DNN warm starts) using 1000 samples for four of the benchmark data sets. All networks have 9 hidden layers with 8 neurons each and ReLU activations. The best method per data set is highlighted in bold
Dataset DEI-MCMC Deep Ensemble
Airfoil $\bm{0.5272}\pm 0.1210$ $-0.1677\pm 0.0766$
Bikesharing $\bm{0.6799}\pm 0.0530$ $0.4145\pm 0.0422$
Concrete $\bm{0.1683}\pm 0.0528$ $-0.2794\pm 0.0622$
Energy $\bm{2.1147}\pm 0.0907$ $0.8019\pm 0.2090$
@@ Appendix D Experimental Settings and Further Details
Our software is implemented in Python and mainly relies on the jax (Bradbury et al., 2018), numpyro (Phan et al., 2019) and pytorch (Paszke et al., 2019) libraries. We further use Docker for a reproducible experimental setup. Our code is available at https://github.com/EmanuelSommer/bnn_connecting_the_dots.
@@
The experiments were run on 4 CPU instances with 32 cores each and 64GB RAM. Sampling 12 chains for most experiments allowed to parallelize the sampling such that at all times 2 experiments with 12 chains each can be run. The sampling with NUTS of 12 chains, 10,000 warmup steps and 8,000 samples (4,000 for the larger data sets) required on average three hours for the smaller data sets, and 30 hours for the larger ones on a two-hidden-layer network with 16 neurons in each hidden layer and tanh activation. This results in about half a second and 8 seconds for 12 parallel samples for the small and large data sets, respectively. Decreasing the number of warmup samples as proposed in DEI-MCMC reduces the number of required samples per chain by more than half, resulting in much faster model fits.